# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v2

@@ B.0.4 Cluster 3: Lonely Moderate Users
B.0.5 Cluster 4: Fulfilled Dependent Users
Fulfilled dependent users (18.32% of sample) show average chatbot usage frequency ($z=0.69$) and high session length ($z=1.76$), with below-average levels of loneliness ($z=-0.56$). This cluster may represent users who feel socially fulfilled by using companion chatbots.
This group demonstrates high extraversion ($z=1.13$), average neuroticism ($z=-0.64$), and high agreeableness ($z=1.21$). They exhibit average levels of socialization ($z=0.42$), and slightly below-average empathy and attraction to a close person. Their trust towards chatbots ($z=0.56$) and expectations from chatbot interactions ($z=0.62$) are above average, with quite high physical attraction towards their chatbot compared to other respondents ($z=1.98$). The group demonstrates above-average perceived impacts on real relationships as a result of using companion chatbots – greater ease and enjoyment in talking to people, but also less engagement and interest in talking to people.
@@ B.0.5 Cluster 4: Fulfilled Dependent Users
B.0.6 Cluster 5: Lonely Light Users
@@
This group demonstrates high neuroticism ($z=1.68$), low conscientiousness ($z=-1.61$), and low self-esteem ($z=-1.25$). Their other characteristics are fairly average, with slightly above-average empathy and attraction towards a close person.
@@
B.0.7 Cluster 6: Socially Challenged Frequent Users
Socially challenged frequent users (10.89% of sample) show average chatbot usage frequency ($z=0.51$) and below-average session length ($z=-0.90$), with high levels of loneliness ($z=1.08$). This cluster represents users who frequently turn to chatbots, possibly due to social challenges.
This group demonstrates low extraversion ($z=-1.30$) and above-average neuroticism ($z=0.86$), along with low levels of socialization ($z=-1.02$), perceived social support ($z=-1.68$), and self-esteem ($z=-1.10$). Their trust towards chatbots ($z=-0.16$) and expectations from chatbot interactions ($z=0.02$) are average. However, trust in familiar and unfamiliar people, and empathy and attraction towards a close person, tend to be very low compared to other users.

### v3

@@ E.0.4 Cluster 3: Lonely Moderate Users
E.0.5 Cluster 4: Fulfilled Dependent Users
Fulfilled dependent users (18.32% of sample) show average chatbot usage frequency ($z=0.69$) and high session length ($z=1.76$), with below-average levels of loneliness (z=0.56z=-0.56). This cluster may represent users who feel socially fulfilled by using companion chatbots.
This group demonstrates high extraversion ($z=1.13$), average neuroticism (z=0.64z=-0.64), and high agreeableness ($z=1.21$). They exhibit average levels of socialization ($z=0.42$), and slightly below-average empathy and attraction to a close person. Their trust towards chatbots ($z=0.56$) and expectations from chatbot interactions ($z=0.62$) are above average, with quite high physical attraction towards their chatbot compared to other respondents ($z=1.98$). The group demonstrates above-average perceived impacts on real relationships as a result of using companion chatbots – greater ease and enjoyment in talking to people, but also less engagement and interest in talking to people.
@@ E.0.5 Cluster 4: Fulfilled Dependent Users
E.0.6 Cluster 5: Lonely Light Users
@@
This group demonstrates high neuroticism ($z=1.68$), low conscientiousness (z=1.61z=-1.61), and low self-esteem (z=1.25z=-1.25). Their other characteristics are fairly average, with slightly above-average empathy and attraction towards a close person.
@@
E.0.7 Cluster 6: Socially Challenged Frequent Users
Socially challenged frequent users (10.89% of sample) show average chatbot usage frequency ($z=0.51$) and below-average session length (z=0.90z=-0.90), with high levels of loneliness ($z=1.08$). This cluster represents users who frequently turn to chatbots, possibly due to social challenges.
This group demonstrates low extraversion (z=1.30z=-1.30) and above-average neuroticism ($z=0.86$), along with low levels of socialization (z=1.02z=-1.02), perceived social support (z=1.68z=-1.68), and self-esteem (z=1.10z=-1.10). Their trust towards chatbots (z=0.16z=-0.16) and expectations from chatbot interactions ($z=0.02$) are average. However, trust in familiar and unfamiliar people, and empathy and attraction towards a close person, tend to be very low compared to other users.